OpenStudy (anonymous):

A rectangular dog pound with three kennels consists of a rectangular fenced area divided by two partitions. Determine the maximum possible area of this pound if 72 yards of chain link fencing is available for its construction.

OpenStudy (anonymous):

First Draw out the figure, you have a rectangle, so Width and Length. We'll choose X to be length, and Y to be width. The area of a rectangle is what? LengthxWidth so A= XY The perimeter is given to us as 72. 2X+2Y = 72. So X+Y = 36 X = 36-y. You can now substitute this back into the area equation. A = (36-Y)*Y It asks for the maximum area, so you'll need to find the derivative and set it = 0. Once you solve the Y value, plug it back in to get X. Remember its asking for area, so you have to Give them X*Y as your final answer

OpenStudy (anonymous):

maximized area is found when x=y, i.e when it is a square. 4 panels for the sides Plus 2 additional for partitions all 6 at equal lengths. 72yds/6 = 12 yards panels. maximun area is 12yd x 12yd = 144 yards^2

OpenStudy (anonymous):